A roulette wheel contains slot numbers 1-36 and a 0 and 00. 18 slots are red and 18 are black. The 0 & 00 are green. Find the probability that the ball lands on red.
Answer by jim_thompson5910(35256)

There are 36+2 = 38 slots and 18 of them are red. So the probability is 18/38 = 9/19